[meteorite-list] Mars Exploration Rovers Update - September 12-20, 2008
From: Ron Baalke <baalke_at_meteoritecentral.com>
Date: Wed, 24 Sep 2008 14:07:48 -0700 (PDT)
SPIRIT UPDATE: Warming Up on Mars - sol 1669-1677, September 12-20, 2008:
With Martian winter on the wane, Spirit is using significantly less
energy to stay warm. During the winter solstice, Spirit needed 90
watt-hours to run the heater. Now, the rover uses between 30 and 40
watt-hours. The reduced demand for power, more than the slow increase in
solar-array input, has freed up energy for other things. In particular,
Spirit has added more images to the 360-degree view of its winter
surroundings, known as the "Bonestell panorama." The top tier, one of
three tiers needed for the final image mosaic, is almost complete.
Plans called for Spirit to use the miniature thermal emission
spectrometer for the first time in several months. The last time the
rover used the instrument was on Martian day, or sol, 1558 (May 21,
2008). On sol 1675 (Sept. 18, 2008), Spirit's schedule of activities
included calibrating the spectometer and using it to observe the sky and
ground. Normally, scientists use the observations to measure
temperatures at different heights and create a temperature profile of
the ground and atmosphere. In this case, the purpose of the measurements
is to verify that the spectrometer is still working after a long, cold
period of disuse. The measurements will also enable scientists to
estimate the amount of dust on the optics. They may or may not provide a
useful temperature profile.
Spirit is healthy and all subsystems are performing as expected as of
the most recent report from NASA's Odyssey orbiter on sol 1674 (Sept.
17, 2008). Solar-array energy has inched upward to 255 watt-hours (100
watt-hours is the amount of energy needed to light a 100-watt bulb for
one hour). Skies are clearer than last week, with tau, a measure of the
amount of sunlight blocked by atmospheric dust, dropping to 0.141.
In addition to taking daily measurements of dust-related changes in
atmospheric opacity (tau), Spirit completed the following activities:
Sol 1669 (Sept. 12, 2008): Spirit recharged the batteries.
Sol 1670: Spirit received new instructions directly from Earth sent at
X-band frequencies to the rover's high-gain antenna. The rover relayed
data to NASA's Odyssey orbiter to be transmitted to Earth.
Sol 1671: Spirit acquired column 20, part 1 and column 21, part 1 of the
full-color Bonestell panorama, using all 13 color filters of the
panoramic camera. The rover took spot images of the sky for calibration
purposes with the panoramic camera.
Sol 1672: Spirit recharged the batteries.
Sol 1673: Spirit recharged the batteries.
Sol 1674: Spirit received new instructions from Earth via the rover's
high-gain antenna. The rover relayed data to Odyssey to be sent to Earth.
Sol 1675: Plans called for Spirit to acquire column 22, part 1 and
column 23, part 1 of the Bonestell panorama and then verify that the
miniature thermal emission spectrometer was still functional. This
involved warming up the actuator, calibrating the instrument, measuring
ground temperature, and measuring atmospheric temperatures at different
heights. Plans also called for Spirit to calibrate the panoramic camera
by taking images in darkness while the instrument was warm.
Sol 1676: Plans called for Spirit to recharge the batteries.
Sol 1677 (Sept. 20, 2008): Plans called for Spirit to recharge the
As of sol 1674 (Sept. 17, 2008), Spirit's total odometry remained at
7,528.0 meters (4.7 miles).
OPPORTUNITY UPDATE: Slipping Like a Dune Buggy - sol 1648-1654,
September 12-18, 2008:
During the past week, Opportunity has been trying to reach a patch of
dust between two crests of the ridge surrounding "Victoria Crater." The
rover approached the ridge from the west, driving on flat ground, on
Martian days, or sols, 1648 and 1650 (Sept. 12 and Sept. 14, 2008).
Then, after reaching a staging position, Opportunity began to climb the
ridge. That's when the rover's wheels began slipping excessively on the
Rover drivers decided to give Opportunity another chance to make it up
the slope by loosening the slip constraints. This allowed Opportunity to
keep trying to climb the slope with a higher rate of wheel slippage. If
the attempt to do this as planned on sol 1654 (Sept. 18, 2008) is not
successful, rover drivers may try a different approach or abandon the
After the dust patch campaign, plans call for Opportunity to drive south
toward a 20-kilometer-wide (12-mile-wide) crater 12 kilometers (7.5
Opportunity is healthy, and all subsystems are performing as expected.
Based on the latest data from sol 1653 (Sept. 17, 2008), the rover has
582 watt-hours of solar power available each day. (One hundred
watt-hours is the amount of energy needed to light a 100-watt bulb for
In addition to measuring dust-related changes in atmospheric clarity
each day with the panoramic camera, Opportunity completed the following
Sol 1648 (Sept. 12, 2008): Opportunity stowed the robotic arm and began
driving toward the dust patch. Just before and after ending the drive,
Opportunity took images with the hazard-avoidance and navigation
cameras, respectively. The rover acquired a 4-by-1 panel of images,
called the "Bagnold mosaic," with the panoramic camera.
Sol 1649: Opportunity took full-color images, using all 13 filters of
the panoramic camera, of a target nicknamed "Drummond." After relaying
data to NASA's Odyssey orbiter for transmission to Earth, Opportunity
measured argon gas in the Martian atmosphere with the alpha-particle
Sol 1650: In the morning, Opportunity took thumbnail images of the sky
with the panoramic camera. The rover continued driving toward the dust
patch and documented progress before and after ending the drive by
taking images with the engineering cameras. Opportunity acquired another
4-by-1 panel of images for the Bagnold mosaic before sending data to
Sol 1651: Opportunity searched for morning clouds in the Martian sky by
taking six time-lapse, movie frames with the navigation camera. The
rover took spot images of the sky for calibration purposes, surveyed the
horizon, and surveyed the sky at low Sun with the panoramic camera.
Sol 1652: In the morning, Opportunity searched for clouds passing
overhead by taking six time-lapse, movie frames with the navigation
camera. The rover checked for drift -- changes with time -- in the
miniature thermal emission spectrometer and also conducted a test of the
instrument. Before beginning the day's drive, Opportunity used the
spectrometer to study a target dubbed "Velvet" and survey the sky and
ground at different elevations. The rover then attempted to drive up the
ridge to the dust patch, acquiring images along the way with the
hazard-avoidance and navigation cameras. Opportunity sent data to
Odyssey for transmission to Earth.
Sol 1653: Opportunity acquired a 3-by-1 mosaic of westward-looking
images with the navigation camera and took images in total darkness with
the panoramic camera for calibration purposes.
Sol 1654 (Sept. 18, 2008): Upon rising, Opportunity took more "dark
current" images with the panoramic camera for calibration purposes. The
rover tried once more to drive to the dust patch, taking images before
and after ending the drive with the hazard-avoidance and navigation
cameras. Before proceeding with plans to measure atmospheric argon,
Opportunity transmitted data to Odyssey for relay to Earth.
As of sol 1653 (Sept. 17, 2008), Opportunity's total odometry was
11,796.22 meters (7.33 miles).
Received on Wed 24 Sep 2008 05:07:48 PM PDT